OpenStudy (anonymous):

Find midpoint AND ABIF a(-3, -5) and B (6, 7)

OpenStudy (fibonaccichick666):

again can you define ABIF and midpoint to begin?

OpenStudy (anonymous):

@FibonacciChick666 http://cs.selu.edu/~rbyrd/math/midpoint/

OpenStudy (fibonaccichick666):

on here please

OpenStudy (anonymous):

\[ \huge \left( \frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2 }\right)\]

OpenStudy (fibonaccichick666):

thanks bat :) so first, what is x1 and y1?

OpenStudy (anonymous):

It's just the average of the two points really. :P

OpenStudy (anonymous):

my teacher said find the ABIF of this problem @iambatman do you know what that means?

OpenStudy (anonymous):

I'm not sure maybe it's in your notes, and he abbreviated something as I'm guessing.

OpenStudy (fibonaccichick666):

I don't know as to what that would stand for...

OpenStudy (anonymous):

Maybe it's asking you to find a kind of line between two points or something, are you sure it just says "ABIF."

OpenStudy (anonymous):

Like find AB is parallel to IF or something...